# Doc2SAR: A Synergistic Framework for High-Fidelity Extraction of Structure-Activity Relationships from Scientific Documents

### v1

@@ Abstract
Extracting molecular structure-activity relationships (SARs) from scientific literature and patents is essential for drug discovery and materials research. However, this task remains challenging due to heterogeneous document formats and limitations of existing methods. Specifically, rule-based approaches relying on rigid templates fail to generalize across diverse document layouts, while general-purpose multimodal large language models (MLLMs) lack sufficient accuracy and reliability for specialized tasks, such as layout detection and optical chemical structure recognition (OCSR). To address these challenges, we introduce DocSAR-200, a rigorously annotated benchmark of 200 scientific documents designed specifically for evaluating SAR extraction methods. Additionally, we propose Doc2SAR, a novel synergistic framework that integrates domain-specific tools with MLLMs enhanced via supervised fine-tuning (SFT). Extensive experiments demonstrate that Doc2SAR achieves state-of-the-art performance across various document types, significantly outperforming leading end-to-end baselines. Specifically, Doc2SAR attains an overall Table Recall of 80.78% on DocSAR-200, exceeding end2end GPT-4o by 51.48%. Furthermore, Doc2SAR demonstrates practical usability through efficient inference and is accompanied by a web app.
@@ 1 Introduction
The extraction of molecular structure-activity relationships (SARs) from scientific literature and patents is fundamental for accelerating the discovery of novel drugs and advanced materials [1, 2, 3]. Accurate and efficient extraction of molecular activity data not only supports computational modeling but also significantly reduces innovation cycles in pharmaceutical and materials science research [4, 5, 6, 7]. Despite its critical importance, automated SAR extraction remains challenging due to the complexity and heterogeneity of scientific documents.
Two major issues currently impede progress in automated SAR extraction. Firstly, there is a notable lack of standardized benchmark datasets specifically designed to evaluate and advance methods for this complex, multimodal information extraction task. Secondly, existing automated extraction methods exhibit critical deficiencies: traditional rule-based methods employing predefined templates are inflexible and fail to generalize across diverse document formats [8, 9, 10]; while general-purpose multimodal large language models (MLLMs), though promising [11, 12], still fall short in specialized subtasks [13, 14, 15, 16]. As our failure analysis in Figure 1 demonstrates, the practical utility of monolithic, end-to-end GPT-4o [17] is severely limited by a cascade of compounding errors. This inherent brittleness highlights the need for a more sophisticated, modular approach.
To address these challenges, we make two primary contributions. First, we introduce DocSAR-200, a meticulously curated benchmark of 200 diverse scientific documents designed to provide a standardized testbed for evaluating SAR extraction methods. Second, to tackle the complexities presented by this benchmark, we propose Doc2SAR, a novel synergistic framework. This framework integrates domain-specific tools like optical chemical structure recognition (OCSR) and fine-tuned MLLMs to precisely solve complex sub-tasks, with robust rule-based methods for post-processing.
Our key contributions can be summarized as follows:
@@
    We introduce DocSAR-200, a comprehensive and rigorously annotated benchmark specifically tailored for evaluating and driving progress in automated SAR extraction.
@@
    Doc2SAR demonstrates strong practical applicability, efficiently processing over 100 PDFs per hour on a single RTX 4090 GPU. Additionally, it features an interactive web interface for easy visualization, verification, and manual refinement, detailed in Appendix E.
@@ 2 DocSAR-200 Benchmark
Figure 2: A representative ground truth annotation from the DocSAR-200 benchmark. For each document, molecular structures and their corresponding entries in activity tables are identified and linked, often across multiple pages, to create a structured record as ground truth in CSV format.
To systematically evaluate and advance automated extraction of SARs from scientific literature, we introduce DocSAR-200, a high-quality benchmark comprising 200 rigorously annotated scientific documents, including both patents and research articles. The benchmark includes complex, real-world cases that require linking molecules and tables across multiple pages, as illustrated by the representative example in Figure 2.
Figure 3: Distribution of SMILES and PDFs across sources. Top row: statistics at the molecules level; bottom row: statistics at the documents level. From left to right: comparison between patent and literature sources; breakdown across literature journals; and distribution across patent authorities.
The final DocSAR-200 dataset includes 3329 valid SMILES, of which 2664 are derived from patent documents and 665 from scientific literature. While patents contribute the majority of molecules, literature-sourced SMILES—primarily from journals such as ACS Med. Chem. Lett. and J. Med. Chem.—tend to exhibit higher semantic clarity and labeling consistency. Overall, the dataset spans 98 patents and 102 research articles, covering documents in English, Chinese, and Japanese, and encompassing a wide range of chemical domains and formatting styles.
Figure 3 provides a structured overview of dataset composition across two axes of analysis. Vertically, the top row aggregates statistics at the molecular (i.e., SMILES) level, while the bottom row reflects the distribution of source documents (i.e., PDFs). Horizontally, each column represents a different source-based perspective: the leftmost column contrasts patent and literature contributions, the middle focuses on literature journal breakdowns, and the rightmost depicts distributions across major patent offices. These statistics provide insight into the heterogeneity and coverage of the benchmark corpus, supporting downstream evaluation under diverse scenarios.
Detailed annotation methodology and quality control procedures are provided in Appendix B.
@@ 3 Doc2SAR Framework
Given an input PDF document $D$, our goal is to extract a set of relationships =$\mathcal{R}=\{(s_{k},a_{k})\}_{k=1}^{N}$, where $s_{k}$ is a molecule’s SMILES string and $a_{k}$ is its associated activity.
To achieve this, we propose Doc2SAR, a novel synergistic framework. Recognizing the inherent brittleness of monolithic, end-to-end MLLMs on this complex task, Doc2SAR employs a modular design that decomposes the problem into independently optimizable components. This architecture achieves superior interpretability and robustness through a synergistic integration of two key elements: specialized, high-precision domain-specific tools for deterministic tasks like OCSR, and powerful, MLLMs for nuanced contextual understanding, such as table extraction and coreference recognition.
As illustrated in Figure 4, the framework operates through three main stages: Layout Detection, Parallel Extraction, and Post-Processing, which are detailed in the following subsections.
@@ 3.1 Layout Detection
Given a raw PDF document $\mathcal{D}$, a YOLO-based [18] layout detection module first segments the document into a set of image blocks, =$\mathcal{B}=\{\mathcal{B}_{\text{mol}},\mathcal{B}_{\text{table}}\}$. Here, $\mathcal{B}_{\text{mol}}$ represents detected molecular images and $\mathcal{B}_{\text{table}}$ represents table regions. This detection function can be formally defined as:
$\mathcal{B}=f_{\text{det}}(\mathcal{D}).$
@@ 3.2 Parallel Extraction
The regions identified by the layout detector are then processed through several concurrent, specialized modules. These modules form the core neural perception engine of our framework.
@@
Each molecular image block $I^{\text{mol}}_{k}\in\mathcal{B}_{\text{mol}}$ is processed by our proprietary OCSR tool. This neural model [19, 20] employs a Swin Transformer [21] as the image encoder $f_{\text{enc}}$ and a BART-style [22] autoregressive decoder $f_{\text{dec}}$ to generate the corresponding SMILES [23] string $s_{k}$. The process is defined as:
$s_{k}=f_{\text{ocsr}}(I^{\text{mol}}_{k})=f_{\text{dec}}(f_{\text{enc}}(I^{\text{mol}}_{k})).$
@@
An MLLM agent extracts the textual identifier for each molecule. For each molecular image $I^{\text{mol}}_{k}$, the agent analyzes the surrounding layout context $C^{\text{ctx}}_{k}$ to identify the corresponding coreference ID $c^{\text{mol}}_{k}$. This operation is defined as:
$c^{\text{mol}}_{k}=f_{\text{coref}}(I^{\text{mol}}_{k},C^{\text{ctx}}_{k}).$
@@
Each table image $I^{\text{table}}_{j}$ is processed by an MLLM agent guided by a conditional prompt. This prompt instructs the model to first screen the table for relevant bioactivity keywords (e.g., $\text{IC}_{50}$). Only if such keywords are found does the model proceed to transform the table image into a structured HTML representation $H_{j}$. This integrated approach simplifies the architecture while leveraging the MLLM’s ability to perform both classification and extraction.
@@ 3.3 Post-processing
As the final stage of the framework, this module synthesizes the outputs from the parallel extraction streams. It performs a crucial cross-modal alignment using the molecular coreference ID as the pivotal key. Specifically, given a molecular pair $(s_{k},c^{\text{mol}}_{k})$ and a tabular entry $(c^{(j)}_{i},a^{(j)}_{i})$, the module employs fuzzy string matching to align their respective IDs. A match is considered valid only if its similarity score exceeds a predefined threshold $\delta$. Each successfully linked pair is then consolidated into a structured SAR record $(s_{k},a^{(j)}_{i})$ for downstream analysis.
3.4 SFT for Key Components
While our framework leverages established tools for Layout Detection and OCSR, the more nuanced tasks of Table HTML Extraction and Molecular Coreference Recognition rely on MLLMs. To overcome the limitations of general-purpose models on these tasks, we employ a SFT strategy to enhance their performance. This involves fine-tuning a single Qwen2.5-VL-7B-Instruct model [24] on a consolidated dataset that combines data for these two sub-tasks. The consolidated training data comprises:
  • Table Extraction Data: 10,000 examples randomly sampled from the PubTabNet [25] training dataset for converting table images to HTML.
  • Molecular Coreference Recognition Data: Our custom dataset of 6,615 instances, including 1,000 synthetic hard cases [26], for identifying molecular identifiers. The detailed construction methodology is provided in Appendix C.
The fine-tuning process was conducted on a multi-GPU server for 3 epochs, using a batch size of 1 and a learning rate of $2\times 10^{-6}$. We adopted rigorous validation protocols throughout the SFT process to ensure the resulting model’s robustness and reliability across both tasks.
@@ 4.1 Experimental Setup
We compare our approach against a suite of powerful, state-of-the-art MLLMs. The baselines include OpenAI’s GPT-4o [17], Google’s Gemini-2.5-flash [27], Meta’s Llama3.2-VL-90B-Instruct [28], and Alibaba’s Qwen2.5-VL [24] series, specifically the 72B-Instruct and 7B-Instruct versions. All models are evaluated in a zero-shot setting using carefully crafted prompts.
@@
For our experiments, we accessed proprietary models like GPT-4o and Gemini-2.5-flash via their official APIs, while open-source models, including the Llama and Qwen series, were hosted locally from public repositories. To ensure a fair and robust comparison, the task-specific prompts used for all models were developed and optimized on a small, disjoint set of validation documents not present in any of the test sets. To promote reproducibility, the DocSAR-200 benchmark and all evaluation scripts will be made publicly available at https://github.com/Jesse-zjx/Doc2SAR.
4.2 Main Results: End-to-End vs. Doc2SAR Framework
This section presents the comprehensive evaluation of our framework on the DocSAR-200 benchmark, with detailed results shown in Table 1. The analysis is structured to first cover the overall performance, followed by granular breakdowns by document type and sub-task.
Table 1: Table Recall (%) on the DocSAR-200 benchmark.
Model Table Recall (%)
Overall Doc. Type Sub. Task
Patents Liters. Coref. Rec. Table. Ex.
# End to End Method
GPT-4o 29.30 20.75 39.19 0.00 49.23
Gemini-2.5-flash 23.03 12.32 40.27 0.00 38.54
Llama3.2-VL-90B-Ins 4.14 0.00 23.62 0.00 7.23
Qwen2.5-VL-72B-Ins 28.51 16.65 45.61 0.00 47.77
# Doc2SAR Framework
GPT-4o 67.68 59.58 79.88 60.58 75.69
Llama3.2-VL-90B-Ins. 64.32 59.57 69.92 55.37 72.18
Qwen2.5-VL-72B-Ins. 75.55 70.67 82.52 64.07 81.03
Qwen2.5-VL-7B-Ins. 73.76 68.27 81.95 59.03 82.62
Ours 80.78 74.09 90.99 65.33 90.08
Overall Performance
The results in Table 1 reveal two key findings regarding overall performance. First, a stark performance gap exists between our synergistic framework and end-to-end methods. End-to-end models struggle significantly, with even powerful models like GPT-4o achieving only 29.30% table recall , and Llama3.2-VL-90B-Ins yielding almost no rational results (4.14%). In contrast, integrating these models into our Doc2SAR framework, which uses domain-specific tools for Layout Detection and OCSR, yields dramatic performance boosts. Second, targeted fine-tuning is critical for state-of-the-art performance. Our SFT-enhanced model (Ours) achieves the highest overall table recall of 80.78%. Notably, our model, fine-tuned on the Qwen2.5-VL-7B-Ins., significantly outperforms the much larger model Qwen2.5-VL-72B-Ins. (80.78% vs. 75.55%).
@@
Analyzing the results by document type shows our framework’s robustness across different formats. The framework approach is particularly effective for patents, which often feature dense text and complex table structures. Furthermore, the SFT process delivers a pronounced improvement on literature articles, raising the recall from 81.95% (for Qwen2.5-VL-7B-Ins) to 90.99% (Ours). This is likely because the PubTabNet data used for fine-tuning aligns well with the typical table layouts found in scientific literature.
@@
The sub-task breakdown uncovers the fundamental weaknesses of the end-to-end approach. All end-to-end models completely fail on the Coreference Recognition sub-task, scoring 0.00% table recall. This failure stems from the inherent inability of current general-purpose MLLMs to perform accurate OCSR and subsequently link the recognized structures to their correct coreferences. Our synergistic framework, by incorporating a dedicated OCSR tool, effectively solves this bottleneck. A similar performance gap is also observed in the Table Extraction sub-task, further validating the necessity of a modular, specialized approach.
4.3 Key Components Analysis: The impact of SFT
To precisely quantify the benefits of SFT, we conducted a granular analysis on the two key MLLM-driven components: Coreference Recognition and Table Structure Recognition. These experiments isolate the impact of SFT by comparing our fine-tuned model against its base architecture and the zero-shot GPT-4o baseline, demonstrating how targeted training enhances capabilities that are critical to our framework’s overall success. The detailed results are presented in Table 2 and Table 3.

### v2

@@ Abstract
Extracting molecular structure-activity relationships (SARs) from scientific literature and patents is essential for drug discovery and materials research. However, this task remains challenging due to heterogeneous document formats and limitations of existing methods. Specifically, rule-based approaches relying on rigid templates fail to generalize across diverse document layouts, while general-purpose multimodal large language models (MLLMs) lack sufficient accuracy and reliability for specialized tasks, such as layout detection and optical chemical structure recognition (OCSR). To address these challenges, we introduce DocSAR-200, a rigorously annotated benchmark of 200 scientific documents designed specifically for evaluating SAR extraction methods. Additionally, we propose Doc2SAR, a novel synergistic framework that integrates domain-specific tools with MLLMs enhanced via supervised fine-tuning (SFT). Extensive experiments demonstrate that Doc2SAR achieves state-of-the-art performance across various document types, significantly outperforming leading end-to-end baselines. Specifically, Doc2SAR attains an overall Table Recall of 80.78% on DocSAR-200, exceeding end2end GPT-4o by 51.48%. Furthermore, Doc2SAR demonstrates practical usability through efficient inference and is accompanied by a web app. The code and data are provided in the supplementary materials.
@@ 1 Introduction
The extraction of molecular structure-activity relationships (SARs) from scientific literature and patents is fundamental for accelerating the discovery of novel drugs and advanced materials (Chu and Fernandes 1989; Farhadi et al. 2019; Hansch 1969). Accurate and efficient extraction of molecular activity data not only supports computational modeling but also significantly reduces innovation cycles in pharmaceutical and materials science research (Drews 2000; Dara et al. 2022; Deng et al. 2022; Sadybekov and Katritch 2023). Despite its critical importance, automated SAR extraction remains challenging due to the complexity and heterogeneity of scientific documents.
Two major issues currently impede progress in automated SAR extraction. Firstly, there is a notable lack of standardized benchmark datasets specifically designed to evaluate and advance methods for this complex, multimodal information extraction task. Secondly, existing automated extraction methods exhibit critical deficiencies: traditional rule-based methods employing predefined templates are inflexible and fail to generalize across diverse document formats (Tharatipyakul et al. 2012; Ibison et al. 1993; Swain and Cole 2016); while general-purpose multimodal large language models (MLLMs), though promising (Zheng et al. 2023; Smith et al. 2025), still fall short in specialized subtasks (Cai et al. 2024; Wadden et al. 2024; Taylor et al. 2022; Beltagy, Lo, and Cohan 2019). As our failure analysis in Figure 1 demonstrates, the practical utility of monolithic, end-to-end GPT-4o (Hurst et al. 2024) is limited by a cascade of compounding errors. This inherent brittleness highlights the need for a more sophisticated, modular approach.
To address these challenges, we make two primary contributions. First, we introduce DocSAR-200, a meticulously curated benchmark of 200 diverse scientific documents designed to provide a standardized testbed for evaluating SAR extraction methods. Second, to tackle the complexities presented by this benchmark, we propose Doc2SAR, a novel synergistic framework. This framework integrates domain-specific tools like optical chemical structure recognition (OCSR) and fine-tuned MLLMs to precisely solve complex sub-tasks, with robust rule-based methods for post-processing. Our key contributions can be summarized as:
@@
    We introduce DocSAR-200, a comprehensive and rigorously annotated benchmark specifically tailored for evaluating progress in automated SAR extraction.
@@
    Doc2SAR demonstrates strong practical applicability, efficiently processing over 100 PDFs per hour on a single RTX 4090 GPU. Additionally, it features an interactive web interface for visualization and manual refinement.
Figure 2: A representative ground truth annotation from the DocSAR-200 benchmark, showing molecular structures linked to activity table entries across multiple pages, formatted as structured CSV records.
@@ 2 DocSAR-200 Benchmark
To systematically evaluate and advance automated extraction of structure–activity relationships (SARs) from heterogeneous scientific sources, we introduce DocSAR-200, a carefully curated benchmark comprising 200 documents (98 patents and 102 research articles) in multiple languages. The design emphasizes three interdependent, hierarchical challenges that any robust extraction system must handle: source diversity, sparse signal localization, and molecular complexity. A concrete example of cross-page molecule-to-table linking is shown in Figure 2.
Figure 3: Overview of DocSAR-200. (a) File-level composition broken down by patents (blue) and literature (red), with further subdivision across journals and patent offices. (b) Distribution of activity table types within 2617 tables; the gray region indicates the proportion of irrelevant (non-activity) tables. (c) Molecular size distribution measured by atom count, including typical and very large molecules.
Figure 3 summarizes the benchmark along three core axes. In Figure 3(a), we expose the breadth and heterogeneity of the source corpus: documents span both patents and peer-reviewed literature, originate from multiple patent authorities and journals, and include multilingual content. This multi-type, multi-source coverage reduces overfitting to a single style or domain and reflects real-world variability in formatting and terminology that extraction systems must tolerate; dataset heterogeneity has been shown to be crucial for robust generalization in information extraction benchmarks and downstream models (Schwartz et al. 2020; Dodge et al. 2021; Gao, Fisch, and Chen 2021). Literature tends to offer clearer semantic signals and terminology consistency, whereas patents contribute scale and diverse phrasing, creating complementary challenges for parsers.
In Figure 3(b), we highlight the difficulty of sparse signal localization. Of the 2617 tables in the benchmark, only a subset encode meaningful activity measurements (e.g., IC50, Ki, Kd), while 77.1% are irrelevant to activity, yielding a “needle-in-a-haystack” scenario where a system must both distinguish relevant from spurious tabular content and correctly interpret a variety of activity metrics. Extracting sparse, meaningful structured data from noisy or abundant background content is a well-known challenge in scientific and biomedical information extraction, addressed in prior work on distant supervision and table understanding (Mintz et al. 2009; Zhong et al. 2019; Li et al. 2022b).
Figure 3(c) characterizes molecular complexity via the distribution of atom counts across benchmark molecules. The histogram shows a concentration around roughly 30 atoms, with a right-skewed tail extending to substantially larger molecules (including some exceeding 100 atoms). The overlaid kernel density estimate (KDE) provides a smoothed approximation of the empirical distribution, making subtle features (such as skew and potential multi-modality) more visible than a raw histogram alone. The divergence between the mean and median further signals right skew, indicating that a minority of large molecules inflates the average size. KDE is a standard nonparametric technique for continuous distribution estimation that avoids restrictive assumptions and can reveal long tails or structure in molecular property distributions (Silverman 1986; Hastie, Tibshirani, and Friedman 2009; Chen, Genovese, and Wasserman 2016). These size variations require molecular recognition components to be robust across a wide dynamic range, a known consideration in AI-enabled drug discovery where molecular size and complexity influence representation fidelity and predictive performance (Vamathevan et al. 2019; Zhou et al. 2020).
Collectively, the three dimensions force extraction pipelines to reconcile heterogeneous source formats, precisely locate sparse but critical activity signals, and handle chemically diverse entities of varying complexity. Detailed annotation methodology and quality control procedures are provided in Appendix B.
@@ 3 Doc2SAR Framework
Given an input PDF document $D$, our goal is to extract a set of relationships $\mathcal{R}=\{(s_{k},a_{k})\}_{k=1}^{N}$, where $s_{k}$ is a molecule’s SMILES string and $a_{k}$ is its associated activity. To achieve this, we propose Doc2SAR, a novel synergistic framework that decomposes the complex SAR extraction task into independently optimizable components. Recognizing the inherent brittleness of monolithic end-to-end MLLMs, Doc2SAR strategically integrates specialized domain-specific tools for deterministic tasks (e.g., OCSR) with fine-tuned MLLMs for contextual understanding (e.g., coreference recognition). As illustrated in Figure 4, the framework operates through three main stages: Layout Detection, Parallel Extraction, and Post-Processing.
@@ 3.1 Layout Detection
We employ a YOLO-based object detection model (Khanam and Hussain 2024) to identify and localize key structural elements within PDF documents. The module converts each PDF page into high-resolution images and extracts spatial coordinates of document elements. For SAR extraction, we focus on two critical categories: molecular structures ($\mathcal{B}_{\text{mol}}$) and tabular regions ($\mathcal{B}_{\text{table}}$). Each detected element is characterized by bounding box coordinates $(x,y,w,h)$ and confidence scores. The detection function is defined as:
$\mathcal{B}=f_{\text{det}}(\mathcal{D})=\{\mathcal{B}_{\text{mol}},\mathcal{B}_{\text{table}}\}$
@@ 3.2 Parallel Extraction
The detected regions are processed through specialized modules that form the core neural perception engine of our framework.
@@
Each molecular image $I^{\text{mol}}_{k}\in\mathcal{B}_{\text{mol}}$ is processed by our specialized OCSR model to generate the corresponding SMILES (Weininger 1988) string $s_{k}$. Our architecture consists of three components: (1) a Swin Transformer (Liu et al. 2021) image encoder $f_{\text{enc}}$ for hierarchical visual feature extraction, (2) a two-layer MLP vision-language connector $f_{\text{conn}}$ following LLaVA (Liu et al. 2023) design, and (3) a BART-style (Lewis et al. 2019) autoregressive decoder $f_{\text{dec}}$ for SMILES generation. The process is defined as:
$s_{k}=f_{\text{ocsr}}(I^{\text{mol}}_{k})=f_{\text{dec}}(f_{\text{conn}}(f_{\text{enc}}(I^{\text{mol}}_{k})))$
To ensure domain-specific accuracy, we fine-tune the model initialized from MolParser (Fang et al. 2024) on 515 manually curated molecular image-SMILES pairs from patents and literature for 20 epochs with learning rate $5\times 10^{-5}$.
@@
Establishing correspondence between molecular structure images and their textual identifiers is critical for SAR extraction. For each molecular image $I^{\text{mol}}_{k}$, an MLLM analyzes surrounding layout context $C^{\text{ctx}}_{k}$ to extract coreference ID $c^{\text{mol}}_{k}$. The context encompasses a spatial window (1.5× original dimensions) including captions, nearby text, table content, and figure annotations. The MLLM processes both image and textual context simultaneously, guided by task-specific prompts to identify compound identifiers such as numbers (e.g., ”1a”, ”compound 5”) or chemical names. The process handles challenging scenarios including molecules in complex figures and structures with multiple identifiers. This operation is defined as:
$c^{\text{mol}}_{k}=f_{\text{coref}}(I^{\text{mol}}_{k},C^{\text{ctx}}_{k};\theta_{\text{prompt}})$
where $\theta_{\text{prompt}}$ represents prompt parameters guiding identifier extraction.
@@
Scientific documents contain numerous tables, but only a subset contains relevant bioactivity data. We implement a two-stage conditional strategy: first screening for bioactivity keywords ($\text{IC}_{50}$, $\text{EC}_{50}$, Ki, Kd), then converting relevant tables to structured HTML. The MLLM handles complex structures including merged cells and multi-level headers while preserving semantic relationships between molecular identifiers and activity values. To ensure quality, the prompt incorporates formatting instructions for numerical precision, cell alignment, and missing data handling. The process is formalized as:
$H_{j}=\begin{cases}f_{\text{html}}(I^{\text{table}}_{j};\theta_{\text{convert}})&\text{if }f_{\text{screen}}(I^{\text{table}}_{j};\theta_{\text{screen}})=\text{True}\\ \emptyset&\text{otherwise}\end{cases}$
@@ 3.3 Post-processing
The post-processing module synthesizes parallel extraction outputs through cross-modal alignment using coreference IDs. The alignment operates on molecular data $\{(s_{k},c^{\text{mol}}_{k})\}_{k=1}^{M}$ and tabular data $\{(c^{(j)}_{i},a^{(j)}_{i})\}_{i=1}^{N_{j}}$ from each table $j$. To handle noise and variability in extracted identifiers, we implement multi-tier fuzzy matching based on normalized Levenshtein distance:
$\text{sim}(c^{\text{mol}}_{k},c^{(j)}_{i})=1-\frac{\text{Levenshtein}(c^{\text{mol}}_{k},c^{(j)}_{i})}{\max(|c^{\text{mol}}_{k}|,|c^{(j)}_{i}|)}$.
The matching employs cascading criteria: exact matching, case-insensitive matching, normalized matching (removing delimiters), and fuzzy matching. A match is valid if similarity exceeds $\delta=0.8$, determined empirically. For multiple candidates, we use confidence-based ranking considering similarity scores, spatial proximity, and naming consistency. Quality control includes SMILES validation, activity range checks, and duplicate removal. Successfully linked pairs form structured SAR records:
$\mathcal{R}=\{(s_{k},a^{(j)}_{i}):\max_{j}\max_{i}\text{sim}(c^{\text{mol}}_{k},c^{(j)}_{i})\geq\delta\}.$
3.4 Supervised Fine-Tuning
While our framework leverages established tools for Layout Detection and OCSR, Table HTML Extraction and Molecular Coreference Recognition require enhanced MLLM capabilities. We employ targeted SFT using a single Qwen2.5-VL-7B-Instruct model (Bai et al. 2025) on consolidated data combining both sub-tasks. Our consolidated training data comprises:
  • Table Extraction Data: 10,000 examples randomly sampled from PubTabNet (Zhong, ShafieiBavani, and Jimeno Yepes 2020), focusing on scientific tables with complex structures similar to those encountered in chemical literature.
  • Molecular Coreference Recognition Data: 6,373 instances, including 5,552 challenging real-world examples extracted from patents and literature and 821 simpler synthetic cases designed to provide coverage and consistency across less ambiguous layouts. The detailed construction methodology for the synthetic portion is provided in Appendix C.
The fine-tuning employs mixed-batch training with examples from both tasks randomly interleaved to promote balanced learning. We train for 3 epochs using batch size 1 and learning rate $2\times 10^{-6}$.
Table 1: Performance comparison on the DocSAR-200 benchmark, measured by Table Recall (%).
Model Table Recall (%)
Overall Doc. Type Sub. Task
Patents Literatures Coref. Rec. Table. Ex.
# End to End Method
GPT-4o (Hurst et al. 2024) 29.30 20.75 39.19 0.00 49.23
Gemini-2.5-flash (Team et al. 2023) 23.03 12.32 40.27 0.00 38.54
Llama3.2-VL-90B-Ins. (Grattafiori et al. 2024) 4.14 0.00 23.62 0.00 7.23
Qwen2.5-VL-72B-Ins. (Bai et al. 2025) 28.51 16.65 45.61 0.00 47.77
# Doc2SAR Framework
GPT-4o (Hurst et al. 2024) 67.68 59.58 79.88 60.58 75.69
Gemini-2.5-flash (Team et al. 2023) 70.32 64.29 78.71 58.94 78.34
Llama3.2-VL-90B-Ins. (Grattafiori et al. 2024) 64.32 59.57 69.92 55.37 72.18
Qwen2.5-VL-72B-Ins. (Bai et al. 2025) 75.55 70.67 82.52 64.07 81.03
Qwen2.5-VL-7B-Ins. (Bai et al. 2025) 73.76 68.27 81.95 59.03 82.62
Ours 80.78 74.09 90.99 65.33 90.08
@@ 4.1 Experimental Setup
We compare our approach against a suite of powerful, state-of-the-art MLLMs. The baselines include OpenAI’s GPT-4o (Hurst et al. 2024), Google’s Gemini-2.5-flash (Team et al. 2023), Meta’s Llama3.2-VL-90B-Instruct (Grattafiori et al. 2024), and Alibaba’s Qwen2.5-VL (Bai et al. 2025) series, specifically the 72B-Instruct and 7B-Instruct versions. All models are evaluated in a zero-shot setting using carefully crafted prompts.
@@
For our experiments, we accessed proprietary models like GPT-4o and Gemini-2.5-flash via their official APIs, while open-source models, including the Llama and Qwen series, were hosted locally from public repositories. To ensure a fair and robust comparison, the task-specific prompts used for all models were developed and optimized on a small, disjoint set of validation documents not present in any of the test sets.
4.2 Main Results: End-to-End vs. Doc2SAR
This section presents our comprehensive evaluation on the DocSAR-200 benchmark (Table 1), demonstrating the superiority of our modular framework over end-to-end approaches. Our analysis reveals three key insights: (1) dramatic performance gaps between methodological approaches, (2) framework robustness across diverse document types, and (3) the critical importance of specialized components for complex sub-tasks.
Overall Performance.
Table 1 reveals a substantial performance gap between our synergistic framework and end-to-end methods. End-to-end approaches show severe limitations: GPT-4o achieves only 29.30% table recall, while Llama3.2-VL-90B struggles at 4.14%. In stark contrast, our Doc2SAR framework transforms these same models, with our SFT-enhanced approach achieving 80.78% table recall, representing a 175% relative improvement over the strongest end-to-end baseline. Remarkably, our fine-tuned 7B model substantially outperforms the much larger 72B variant (80.78% vs. 75.55%), demonstrating that domain-specific training trumps raw model scale for specialized scientific document understanding tasks.
@@
Our framework demonstrates robust performance across diverse document formats, with particularly strong results on literature articles (90.99%) compared to patents (74.09%). This gap likely reflects structural differences: literature articles feature more standardized table layouts that align well with our PubTabNet training data, while patents exhibit greater format diversity, dense technical language, and complex nested structures that challenge automated processing. The SFT process delivers substantial improvements for literature (81.95% → 90.99%), representing an 11% relative gain that validates our targeted training strategy. The challenging patents highlight the critical value of our synergistic approach, where specialized OCSR and MLLM components collaborate to handle complex, heterogeneous layouts.
@@
The sub-task breakdown exposes fundamental limitations of end-to-end approaches. Most critically, all end-to-end models achieve 0.00% on Coreference Recognition, revealing their inability to perform accurate OCSR followed by cross-modal alignment (a complex multi-step process requiring both precise molecular recognition and contextual reasoning capabilities essential for SAR extraction). Our modular design addresses this bottleneck through dedicated OCSR tools, achieving 65.33% on this challenging task. Similarly, for Table Extraction, our approach (90.08%) dramatically outperforms the best end-to-end method (49.23%), with an 83% relative improvement that underscores the value of specialized table understanding over generic multimodal processing, particularly for scientific tables with domain-specific terminology and complex structures. These results validate our core hypothesis that complex scientific document understanding requires modular, domain-specific architectures rather than monolithic end-to-end solutions.
4.3 Components Analysis: The impact of SFT
To quantify the benefits of our targeted SFT strategy, we conduct isolated evaluations of the two key MLLM-driven components: Molecular Coreference Recognition and Table HTML Extraction. These experiments compare our fine-tuned model against Qwen2.5-VL-7B-Instruct and GPT-4o baseline, demonstrating how domain-specific training enhances capabilities critical to our framework’s success. Results are presented in Tables 2 and 3.